OpenStudy (anonymous):

Assuming that the smallest measurable wavelength in an experiment is 0.890 fm (femtometers), what is the maximum mass of an object traveling at 563 m·s–1 for which the de Broglie wavelength is observable?

OpenStudy (aaronq):

1 fm = \(1.0*10^{-15}\) m so, \(\lambda=\dfrac{h}{mv}\rightarrow 0.890*10^{-15} m=\dfrac{6.626*10^{-34}m^2kg/s}{m(563 \;m/s)}\) \(mass=\dfrac{(6.626*10^{-34}m^2kg/s)}{(563 \;m·s^{–1})*(0.890*10^{-15} m)}\)